# Non-Hermitian Linear Electro-Optic Effect Through Interactions of Free and Bound Charges

### v1

@@ IV Spatial symmetries
Next, we discuss the spatial symmetries compatible with nontrivial nonlinear couplings and derive explicitly the linearized permittivity for materials belonging to the 2mm symmetry group.
@@ IV.1 Constraints due to spatial symmetries
Among the non-centrosymmetric systems that possess an axis of rotation, it can be checked that rotations by an angle of $\pi$, $32\pi/$ 3, $\pi/2$, and $\pi/3$ are all compatible with the existence of a nontrivial coupling tensor $\mathbf{\overline{C}}_{21}$. For example, we checked that a nontrivial $\mathbf{\overline{C}}_{21}$ is compatible with the symmetries of the 2, 222, 2mm and 32 crystallographic groups (Hermann–Mauguin notation).
IV.2 Linearized permittivity for the 2mm symmetry group
In the rest of the paper, we focus on material belonging to the 2mm symmetry group. Materials in this crystallographic group have two mirror planes and are invariant under a 2-fold rotation symmetry about some specific axis. We note in passing that a MOSFET transistor exhibits the same type of symmetries Lannebère et al. (2022).
We assume, without loss of generality, that the axis of rotation of the structure is parallel to the $z$ direction. The nontrivial elements of the 2mm symmetry group consists of a rotation by an angle of $\pi$ ($\mathcal{R}=\text{diag}\left\{-1,-1,1\right\}$) and two mirror symmetries with respect to planes containing the rotation axis (here taken as $\mathcal{R}=\text{diag}\left\{-1,1,1\right\}$ and $\mathcal{R}=\text{diag}\left\{1,-1,1\right\}$). When these three symmetries are individually imposed to our model, it can be shown that most coupling coefficients in Eq. (19) must vanish. The most general form of the coupling tensor is:
@@ IV.2 Linearized permittivity for the 2mm symmetry group
Next, we calculate the linearized permittivity (25) for a material that exhibits $2mm$-symmetry. Importantly, two cases can be distinguished depending on whether the direction of the drift velocity is parallel or orthogonal to the axis of rotation of the structure. For a drift velocity parallel to the axis of rotation ($\mathbf{v}_{0}=v_{0z}\hat{\mathbf{e}}_{z}$), it can be shown that the linearized permittivity (25) is always a diagonal tensor and therefore describes a reciprocal response. This case will not be further considered in this paper.
@@
$\sum_{j}\left(\omega_{0j}^{2}-\omega^{2}-i\gamma\omega\right)\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\cdot{\mathbf{P}_{\omega}}$ $=\varepsilon_{0}\omega_{b}^{2}\left\{{\mathbf{E}_{\omega}}+qn_{0}v_{0x}\left[(2a_{zxx}\hat{\mathbf{e}}_{z}\otimes\hat{\mathbf{e}}_{x}+\left(a_{xzx}+a_{xxz}\right)\hat{\mathbf{e}}_{x}\otimes\hat{\mathbf{e}}_{z})\cdot{\mathbf{v}_{\omega}}\right.\right.$
@@
This set of equations explicitly shows the form of linearized local field induced by the coupling between the free and bound electrons. As seen, the induced local fields involve the interaction between the $x$ component of $\mathbf{v}_{\omega}$ or $-i\omega\mathbf{P}_{\omega}$ with the $z$ component of $\mathbf{v}_{\omega}$ or $-i\omega\mathbf{P}_{\omega}$, or vice-versa, similar to what happens in standard MOSFET transistors Sze et al. (2021); Lannebère et al. (2022).
Furthermore, it is clear from the above equation that for a drift velocity orthogonal to the material rotation axis, only 5 coefficients out of the 14 coefficients appearing in Eq. (30) play a role in the linearized response. The other coefficients give rise to higher order responses and disappear upon linearization. In fact, from (25) and (27a), one readily finds that for materials belonging to the 2mm symmetry group with a drift orthogonal to the rotation axis, the linearized permittivity tensor is given by:
@@
In the presence of nonlinear coupling between the free and bound electrons, when some of the coefficients $a_{ijk}$ and $b_{ijk}$ are different from zero, the permittivity tensor is non-diagonal. Additionally, due to collisions ($\Gamma>0$ and $\gamma>0$), the functions $\varepsilon_{ii}^{0}$, $f_{1}$ and $g_{ij}$ (with $i,j=x,y,z$) are all complex-valued. Remarkably, all types of coupling result in a permittivity tensor that is asymmetric, leading to a nonreciprocal ($\mathbf{\overline{\varepsilon}}\neq\mathbf{\overline{\varepsilon}}^{T}$) electromagnetic response. In particular, the linearized electro-optic response is not TR symmetric ($\mathbf{\overline{\chi}}_{\rm{f-b}}\neq\mathbf{\overline{\chi}}_{\rm{f-b}}^{\ast}$); note that the tensor $\mathbf{\overline{\chi}}_{\rm{f-b}}$ is formed by the out-of-diagonal elements of $\mathbf{\overline{\varepsilon}}$. This behavior is expected from the presence of a drift current, which is inherently odd under TR symmetry. The broken TR symmetry distinguishes our system from the idealized model described in Ref. Lannebère et al. (2022), aligning it more closely with the behavior observed in actual MOSFET transistors, which also support a DC current Fernandes et al. (2024).
Additionally, all couplings result in a non-Hermitian electro-optic response ($\mathbf{\overline{\chi}}_{\rm{f-b}}\neq\mathbf{\overline{\chi}}_{\rm{f-b}}^{\dagger}$), Furthermore, it can be checked that the tensor ${\mathbf{\overline{\chi}}_{{\rm{NH}}}}=\frac{1}{{2i}}\left({{{\mathbf{\overline{\chi}}}_{{\rm{f-b}}}}-\mathbf{\overline{\chi}}_{{\rm{f-b}}}^{\dagger}}\right)$ is an indefinite matrix. This indicates that the electro-optic interaction may result in either dissipation or in gain Lannebère et al. (2025b). This idea will be further developed in the next section.
Therefore, the nonlinear interactions between the bound and free electrons provide an alternative mechanism to implement the NHEO effect, not directly relying on the Berry curvature dipole. It is interesting to note that the frequency dependence of the tensor $\mathbf{\overline{\chi}}_{\rm{f-b}}$ is qualitatively similar to that of the $\chi^{(2)}$ tensor derived from band theory Boyd (2008) (Sec. 3.2). Thus, in principle it is possible to link the coefficients $a_{ijk}$ and $b_{ijk}$ with the electronic band-structure of the material. Such analysis will not be pursued here, and is left for future work.
@@
In summary, in this section it was demonstrated that a metallic system with the symmetry of the 2mm group, with the static electric bias orthogonal to the rotation axis of the structure, can be used to implement the NHEO effect and achieve nonreciprocity and gain. Different from other platforms based on drifting electrons Borgnia et al. (2015); Duppen et al. (2016); Morgado and Silveirinha (2018); Wenger et al. (2018); Bliokh et al. (2018); Morgado and Silveirinha (2020, 2021); Zhao et al. (2021); Dong et al. (2021), our solution does not rely on Doppler shifts and on waves with extremely short wavelengths (plasmons). Furthermore, in our system the drift velocity does not need to be a significant fraction of the light velocity.
We expect that our model can be implemented either in engineered semiconductor superlattices Rosencher et al. (1996) or in natural materials. For example, potential candidates could be found in the class of the polar/ferroelectrics metals Zhou and Ariando (2020), materials that are both conducting and polarizable. WTe2, a polar metal (as well as a Weyl semimetal) with the symmetry of the 2mm symmetry group could be a possible candidate.
@@ V Wave propagation
To illustrate the properties of our system, next we study the wave propagation in a material described by the linearized permittivity (32). For simplicity, we assume an isotropic Lorentz model such that $\omega_{0j}=\omega_{0}$ for all $j$. Also we focus on a specific type of nonreciprocal and non-Hermitian response and assume that $a_{zxx}$ is the only nonzero coupling coefficient in Eq. (19). Thus, the coupling matrix is $\mathbf{\overline{C}}_{21}=a_{zxx}v_{x}\hat{\mathbf{e}}_{x}\otimes\hat{\mathbf{e}}_{z}$, so that it couples the $x$ and $z$ components of the velocity and polarization current. From Eq. (32), the linearized permittivity satisfies
@@ V.1 Band structure and stability
To begin with, we characterize the photonic band diagrams in the absence of coupling ($\varepsilon_{c}=0$). In this case, the permittivity tensor is isotropic $\mathbf{\overline{\varepsilon}}=\varepsilon_{d}\mathbf{1}_{3\times 3}$ and the plane wave solutions of the homogeneous wave equation ××$\nabla\times\nabla\times\mathbf{E}=\frac{\omega^{2}}{c^{2}}\varepsilon_{d}\mathbf{E}$ can be separated into a longitudinal mode and a doubly degenerated transverse electromagnetic (TEM) mode.
Figure 2: (a) Band diagram of the unbiased material showing the frequency as a function of the normalized wavenumber for $\omega_{b}=0.9\omega_{p}$ and $\omega_{0}=0.3\omega_{p}$. (b) Zoom in near the lower edge of the low frequency band. (c) Zoom in near the lower edge of the high frequency band. The horizontal green solid lines represent the longitudinal mode and the red lines represent the doubly degenerated TEM modes.
@@
When the coupling is nonzero ($\varepsilon_{c}\neq 0$), the material becomes anisotropic. First, we consider wave propagation in the $xoz$ plane, with the wavevector $\mathbf{k}=k\left(\sin(\theta)\hat{\mathbf{e}}_{x}+\cos(\theta)\hat{\mathbf{e}}_{z}\right)$ parameterized by an angle $\theta$ represented in the inset of Fig. 3 (a).
Figure 3: Dispersion of the plane wave modes $\omega=\omega^{\prime}+i\omega^{\prime\prime}$ vs. $\mathbf{k}$ for a real-valued wave vector and propagation in the $xoz$ plane. The medium is described by the permittivity tensor (33). (a) and (b): $\omega^{\prime}$ as a function of $k$ near the edge the low-frequency (high-frequency) branches, respectively, for $\theta=\pi/4$. These two plots are nearly insensitive to the values of $\theta$, $\Gamma$ and $\gamma$. The wavevector in the $xoz$ plane is represented in the inset of (a). (c) and (d): Projection of the band diagram in the complex plane for the low-frequency (high-frequency) bands, respectively, ignoring collisions ($\Gamma=\gamma=0^{+}$). (e) and (f): Similar to (c) and (d) but for $\Gamma=3.85\cdot 10^{-3}\omega_{p}$ and $\gamma=1.232\cdot 10^{-3}\omega_{p}$. For TM waves, the solid red curves in (c)–(f) represent the wave dispersion for $\theta=\pi/4$ and the red shaded regions represent the frequency locus for all angles $\theta$. In all the plots, $\omega_{b}=0.9\omega_{p}$, $\omega_{0}=0.3\omega_{p}$ and $\varepsilon_{0}v_{0x}a_{zxx}=0.01/\omega_{p}$.
In this case, the plane wave solutions of the homogeneous wave equation ××$\nabla\times\nabla\times\mathbf{E}=\frac{\omega^{2}}{c^{2}}\mathbf{\overline{\varepsilon}}\cdot\mathbf{E}$ consist of transverse magnetic (TM) and TEM modes. The TEM mode (ordinary wave) is characterized by an electric field polarized along the $y$ direction and, similar to the unbiased case, by the wavenumber $k=\frac{\omega}{c}\sqrt{\varepsilon_{d}}$. The TM mode (extraordinary wave) is characterized by
@@
where $A_{\text{TM}}$ is a constant and $\nu=\frac{\varepsilon_{d}-\varepsilon_{\text{TM}}\cos(\theta)^{2}}{\varepsilon_{\text{TM}}\cos(\theta)\sin(\theta)-i\varepsilon_{c}}$ is a complex-valued function that depends on $\omega$ and $\theta$. For propagation in the $xoz$ plane, the TEM and TM eigenmodes are orthogonal. It is seen from equation (34b) that even in the absence of collisions (i.e. $\varepsilon_{d}$ and $\varepsilon_{c}$ real), the effective permittivity of the TM mode $\varepsilon_{\text{TM}}$ is complex valued, with an imaginary part that is maximized/minimized for $\theta=\mp\pi/4$ and $\theta=\pm 3\pi/4$. Thus in the absence of collisions, depending on the value of $\theta$, $\varepsilon_{\text{TM}}$ can have either a positive or negative imaginary part implying that the material may either dissipate or amplify an electromagnetic wave. An intriguing consequence is that an antenna embedded in this material would radiate a beam that amplifies within a specific angular region while being suppressed in other directions.

### v2

@@ IV Spatial symmetries
Next, we discuss the spatial symmetries compatible with nontrivial nonlinear couplings and derive explicitly the linearized permittivity for materials belonging to the 2mm point group.
@@ IV.1 Constraints due to spatial symmetries
Among the non-centrosymmetric systems that possess an axis of rotation, it can be checked that rotations by an angle of $\pi$, $2\pi/3$, $\pi/2$, and $\pi/3$ are all compatible with the existence of a nontrivial coupling tensor $\mathbf{\overline{C}}_{21}$. For example, we checked that a nontrivial $\mathbf{\overline{C}}_{21}$ is compatible with the symmetries of the 2, 222, 2mm and 32 crystallographic point groups (Hermann–Mauguin notation).
IV.2 Linearized permittivity for the 2mm point group
In the rest of the paper, we focus on material belonging to the 2mm point group. Materials in this crystallographic point group have two mirror planes and are invariant under a 2-fold rotation symmetry about some specific axis. We note in passing that a MOSFET transistor exhibits the same type of symmetries Lannebère et al. (2022).
We assume, without loss of generality, that the axis of rotation of the structure is parallel to the $z$ direction. The nontrivial elements of the 2mm point group consist of a rotation by an angle of $\pi$ ($\mathcal{R}=\text{diag}\left\{-1,-1,1\right\}$) and two mirror symmetries with respect to planes containing the rotation axis (here taken as $\mathcal{R}=\text{diag}\left\{-1,1,1\right\}$ and $\mathcal{R}=\text{diag}\left\{1,-1,1\right\}$). When these three symmetries are individually imposed to our model, it can be shown that most coupling coefficients in Eq. (19) must vanish. The most general form of the coupling tensor is:
@@ IV.2 Linearized permittivity for the 2mm point group
Next, we calculate the linearized permittivity (25) for a material that exhibits $2mm$ point group symmetry. Importantly, two cases can be distinguished depending on whether the direction of the drift velocity is parallel or orthogonal to the axis of rotation of the structure. For a drift velocity parallel to the axis of rotation ($\mathbf{v}_{0}=v_{0z}\hat{\mathbf{e}}_{z}$), it can be shown that the linearized permittivity (25) is always a diagonal tensor and therefore describes a reciprocal response. This case will not be further considered in this paper.
@@
[…]
@@
This set of equations explicitly shows the form of linearized local field induced by the coupling between the free and bound electrons. As seen, the induced local fields involve the interaction between the $x$ component of $\mathbf{v}_{\omega}$ or $i\omega\mathbf{P}_{\omega}$ with the $z$ component of $\mathbf{v}_{\omega}$ or $i\omega\mathbf{P}_{\omega}$, or vice-versa, similar to what happens in standard MOSFET transistors Sze et al. (2021); Lannebère et al. (2022).
Furthermore, it is clear from the above equation that for a drift velocity orthogonal to the material rotation axis, only 5 coefficients out of the 14 coefficients appearing in Eq. (30) play a role in the linearized response. The other coefficients give rise to higher order responses and disappear upon linearization. In fact, from (25) and (27a), one readily finds that for materials belonging to the 2mm point group with a drift orthogonal to the rotation axis, the linearized permittivity tensor is given by:
@@
In the presence of nonlinear coupling between the free and bound electrons, when some of the coefficients $a_{ijk}$ and $b_{ijk}$ are different from zero, the permittivity tensor is non-diagonal. Additionally, due to collisions ($\Gamma>0$ and $\gamma>0$), the functions $\varepsilon_{ii}^{0}$, $f_{1}$ and $g_{ij}$ (with $i,j=x,y,z$) are all complex-valued. Remarkably, all types of coupling result in a permittivity tensor that is asymmetric, leading to a nonreciprocal ($\mathbf{\overline{\varepsilon}}\neq\mathbf{\overline{\varepsilon}}^{T}$) electromagnetic response. In particular, the linearized electro-optic response is not TR symmetric ($\mathbf{\overline{\chi}}_{\text{f-b}}\neq\mathbf{\overline{\chi}}_{\text{f-b}}^{\ast}$); note that the tensor $\mathbf{\overline{\chi}}_{\text{f-b}}$ is formed by the out-of-diagonal elements of $\mathbf{\overline{\varepsilon}}$. This behavior is expected from the presence of a drift current, which is inherently odd under TR symmetry. The broken TR symmetry distinguishes our system from the idealized model described in Ref. Lannebère et al. (2022), aligning it more closely with the behavior observed in actual MOSFET transistors, which also support a DC current Fernandes et al. (2024).
Additionally, all couplings result in a non-Hermitian electro-optic response ($\mathbf{\overline{\chi}}_{\text{f-b}}\neq\mathbf{\overline{\chi}}_{\text{f-b}}^{\dagger}$), Furthermore, it can be checked that the tensor $\mathbf{\overline{\chi}}_{\text{NH}}=\frac{1}{2i}\left(\mathbf{\overline{\chi}}_{\text{f-b}}-\mathbf{\overline{\chi}}_{\text{f-b}}^{\dagger}\right)$ is an indefinite matrix. This indicates that the electro-optic interaction may result in either dissipation or in gain Lannebère et al. (2025b). This idea will be further developed in the next section.
Therefore, the nonlinear interactions between the bound and free electrons provide an alternative mechanism to implement the NHEO effect, not directly relying on the Berry curvature dipole. It is interesting to note that the frequency dependence of the tensor $\mathbf{\overline{\chi}}_{\text{f-b}}$ is qualitatively similar to that of the $\mathbf{\overline{\chi}}^{(2)}$ tensor derived from band theory Aversa and Sipe (1995); Shen (2003); Boyd (2008) (Sec. 2.2 and 3.2, respectively). Thus, in principle, it is possible to link the coefficients $a_{ijk}$ and $b_{ijk}$ with the electronic band-structure of the material. Even though no data are currently available to enable such a mapping for materials of primary interest, we hope that the present model can serve as a foundation for future theoretical or experimental investigations.
@@
In summary, in this section it was demonstrated that a metallic system with the symmetry of the 2mm point group, with the static electric bias orthogonal to the rotation axis of the structure, can be used to implement the NHEO effect and achieve nonreciprocity and gain, exclusively through the nonlinearity arising from interband interactions not captured by the anomalous velocity. Different from other platforms based on drifting electrons Borgnia et al. (2015); Duppen et al. (2016); Morgado and Silveirinha (2018); Wenger et al. (2018); Bliokh et al. (2018); Morgado and Silveirinha (2020, 2021); Zhao et al. (2021); Dong et al. (2021), our solution does not rely on Doppler shifts and on waves with extremely short wavelengths (plasmons). Furthermore, in our system the drift velocity does not need to be a significant fraction of the light velocity.
We expect that our model can be used to describe the physics of electro-optic effects due to interband interactions (not captured by the semiclassical Boltzmann approach) in low-symmetry metals, in material platforms such as Weyl semimetals Guo et al. (2023), transition metal dichalcogenides (TMDs), chiral perovskites Long et al. (2020) or polar/ferroelectrics metals (materials that are both conducting and polarizable) Zhou and Ariando (2020); Hickox-Young et al. (2023); Bhowal and Spaldin (2023), and engineered semiconductor superlattices Rosencher et al. (1996). A particularly interesting candidate material is WTe2, a TMD that is simultaneously a polar metal and a Weyl semimetal with the symmetry of the 2mm point group.
@@ V Wave propagation
To illustrate the properties of our system, next we study the wave propagation in a material described by the linearized permittivity (32). For simplicity, we assume an isotropic Lorentz model such that $\omega_{0j}=\omega_{0}$ for all $j$. Also we focus on a specific type of nonreciprocal and non-Hermitian response and assume that $a_{zxx}$ is the only nonzero coupling coefficient in Eq. (19). Thus, the coupling matrix is $\mathbf{\overline{C}}_{21}=a_{zxx}v_{x}\hat{\mathbf{e}}_{z}\otimes\hat{\mathbf{e}}_{x}$, so that it couples the $x$ and $z$ components of the velocity and polarization current. From Eq. (32), the linearized permittivity satisfies
@@ V.1 Band structure and stability
To begin with, we characterize the photonic band diagrams in the absence of coupling ($\varepsilon_{c}=0$). In this case, the permittivity tensor is isotropic $\mathbf{\overline{\varepsilon}}=\varepsilon_{d}\mathbf{1}_{3\times 3}$ and the plane wave solutions of the homogeneous wave equation $\nabla\times\nabla\times\mathbf{E}=\frac{\omega^{2}}{c^{2}}\varepsilon_{d}\mathbf{E}$ can be separated into a longitudinal mode and a doubly degenerated transverse electromagnetic (TEM) mode.
Figure 2: (a) Band diagram of the unbiased material showing the frequency as a function of the normalized wavenumber for $\omega_{\text{b}}=0.9\omega_{p}$ and $\omega_{0}=0.3\omega_{p}$. (b) Zoom in near the lower edge of the low frequency band. (c) Zoom in near the lower edge of the high frequency band. The horizontal green solid lines represent the longitudinal mode and the red lines represent the doubly degenerated TEM modes.
@@
When the coupling is nonzero ($\varepsilon_{c}\neq 0$), the material becomes anisotropic. First, we consider wave propagation in the xozxoz plane, with the wavevector $\mathbf{k}=k\left(\sin(\theta)\hat{\mathbf{e}}_{x}+\cos(\theta)\hat{\mathbf{e}}_{z}\right)$ parameterized by an angle $\theta$ represented in the inset of Fig. 3 (a).
Figure 3: Dispersion of the plane wave modes $\omega=\omega^{\prime}+i\omega^{\prime\prime}$ vs. $\mathbf{k}$ for a real-valued wave vector and propagation in the xozxoz plane. The medium is described by the permittivity tensor (33). (a) and (b): $\omega^{\prime}$ as a function of $k$ near the edge the low-frequency (high-frequency) branches, respectively, for $\theta=\pi/4$. These two plots are nearly insensitive to the values of $\theta$, $\Gamma$ and $\gamma$. The wavevector in the xozxoz plane is represented in the inset of (a). (c) and (d): Projection of the band diagram in the complex plane for the low-frequency (high-frequency) bands, respectively, ignoring collisions ($\Gamma=\gamma=0^{+}$). (e) and (f): Similar to (c) and (d) but for $\Gamma=3.85\cdot 10^{-3}\omega_{p}$ and $\gamma=1.232\cdot 10^{-3}\omega_{p}$. For TM waves, the solid red curves in (c)–(f) represent the wave dispersion for $\theta=\pi/4$ and the red shaded regions represent the frequency locus for all angles $\theta$. In all the plots, $\omega_{\text{b}}=0.9\omega_{p}$, $\omega_{0}=0.3\omega_{p}$ and $\varepsilon_{0}v_{0x}a_{zxx}=0.01/\omega_{p}$.
In this case, the plane wave solutions of the homogeneous wave equation $\nabla\times\nabla\times\mathbf{E}=\frac{\omega^{2}}{c^{2}}\mathbf{\overline{\varepsilon}}\cdot\mathbf{E}$ consist of transverse magnetic (TM) and TEM modes. The TEM mode (ordinary wave) is characterized by an electric field polarized along the $y$ direction and, similar to the unbiased case, by the wavenumber $k=\frac{\omega}{c}\sqrt{\varepsilon_{d}}$. The TM mode (extraordinary wave) is characterized by
@@
where $A_{\text{TM}}$ is a constant and $\nu=\frac{\varepsilon_{d}-\varepsilon_{\text{TM}}\cos(\theta)^{2}}{\varepsilon_{\text{TM}}\cos(\theta)\sin(\theta)-i\varepsilon_{c}}$ is a complex-valued function that depends on $\omega$ and $\theta$. For propagation in the xozxoz plane, the TEM and TM eigenmodes are orthogonal. It is seen from equation (34b) that even in the absence of collisions (i.e. $\varepsilon_{d}$ and $\varepsilon_{c}$ real), the effective permittivity of the TM mode $\varepsilon_{\text{TM}}$ is complex valued, with an imaginary part that is maximized/minimized for $\theta=\mp\pi/4$ and $\theta=\pm 3\pi/4$. Thus in the absence of collisions, depending on the value of $\theta$, $\varepsilon_{\text{TM}}$ can have either a positive or negative imaginary part implying that the material may either dissipate or amplify an electromagnetic wave. An intriguing consequence is that an antenna embedded in this material would radiate a beam that amplifies within a specific angular region while being suppressed in other directions.